OpenStudy (anonymous):

1/2+4/2x=x+4/10 solve for x

OpenStudy (allank):

First, get rid of the denominator on the right side of the equal sign by multiplying through by 10 to get: 5+20x=x+4 then simplify by adding the like terms and computing the value of x.

hero (hero):

\(\large\frac{1}{2} + \frac{4}{2x} = \frac{x+4}{10} \) Multiply the first fraction by x/x : \(\large\frac{x}{2x} + \frac{4}{2x} = \frac{x+4}{10} \) Combine the fractions on the left side to get \(\large\frac{x+4}{2x} = \frac{x+4}{10} \) Cross multiply to get: 2x(x+4) = 10(x+4) \(2x^2 + 8x = 10x + 40 \\ 2x^2 - 2x - 40 = 0 \\ 2(x^2 - x - 20) = 0 \\ x^2 - x - 20 = 0 \\ x^2 - 5x + 4x - 20 = 0 \\ x(x-5) + 4(x-5) = 0 \\ (x-5)(x+4) = 0 \\ x = 5, x = -4\)

OpenStudy (allank):

Nice.